# On the convergence of loss and uncertainty-based active learning algorithms

### v2

@@ Abstract
We consider the convergence rates of loss and uncertainty-based active learning algorithms under various assumptions. Firstly, we establish a set of conditions that ensure convergence rates when applied to linear classifiers and linearly separable datasets. This includes demonstrating convergence rate guarantees for loss-based sampling with various loss functions. Secondly, we introduce a framework that allows us to derive convergence rate bounds for loss-based sampling by leveraging known convergence rate bounds for stochastic gradient descent algorithms. Lastly, we propose a new algorithm that combines point sampling and stochastic Polyak’s step size. We establish a condition on the sampling process, ensuring a convergence rate guarantee for this algorithm, particularly in the case of smooth convex loss functions. Our numerical results showcase the efficiency of the proposed algorithm.
@@ 1 Introduction
In practice, when training machine learning models for prediction tasks (classification or regression), one often has an abundance of unlabeled data available, but obtaining the corresponding labels may come at a high cost. This may especially be the case in fields like computer vision, natural language processing, and speech recognition. Active learning algorithms are designed to efficiently learn a prediction model by employing a label acquisition, with the goal of minimizing the number of labels used to train an accurate prediction model.
@@
Various label acquisition strategies have been proposed, each aiming to select informative points for the underlying model training task. For instance, active learning strategies include query-by-committee (Seung et al., 1992), expected model change (Settles et al., 2007), expected error reduction (Roy and McCallum, 2001), expected variance reduction (Wang et al., 2016), and mutual information maximization (Kirsch et al., 2019; Kirsch and Gal, 2022).
@@
A common label acquisition strategy involves estimating uncertainty, which can be viewed as self-disagreement about predictions made by a given model. Algorithms using an uncertainty acquisition strategy are referred to as uncertainty-based active learning algorithms. Recently, a loss-based active learning approach gained attention in research Yoo and Kweon (2019), Lahlou et al. (2022), Nguyen et al. (2021), Luo et al. (2021), and is now applied at scale in industry, such as for training integrity violation classifiers at Meta. This method involves selecting points for which there is a disagreement between the prediction model and an oracle, as measured by a loss function. Since the true loss of a data point is unknown prior to the acquisition of the label, in practice it is estimated using supervised learning. Loss-based sampling aligns with the spirit of the perceptron algorithm (Rosenblatt, 1958), which updates the model only for falsely-classified points.
@@
Convergence guarantees for some uncertainty sampling strategies have recently been established, such as for margin of confidence sampling (Raj and Bach, 2022). As for loss-based active learning algorithms, there are only limited results on their convergence properties, as they have been started to be studied only recently. e.g. Liu and Li (2023).
The primary focus of this paper is to establish convergence guarantees for loss and uncertainty-based learning algorithms. We delve into stream-based algorithms, where model training entails the use of a stochastic gradient descent algorithm with a loss or an uncertainty-based acquisition function. Our work provides new results on conditions that ensure the convergence of loss- and uncertainty-based active learning algorithms, along with bounds on convergence rates, accommodating various loss functions.
@@
In the case of loss-based acquisition strategies, we assume that the active learner has access to an oracle providing an unbiased estimate of the conditional expected loss for a point, given the feature vector of the point and the current prediction model. In practical settings, the loss cannot be evaluated at acquisition time since labels are yet unknown. Instead, a separate prediction model can be devised, which can estimate the loss subject to estimation noise. In our experiments, we assess the impact of the noise in such a loss estimator for various loss-based active learning strategies.
@@
It is noteworthy that uncertainty and loss-based acquisition strategies are also of interest for the data subset selection problem, often referred to as core-set selection or data pruning. This problem involves finding a small subset of training data such that the predictive performance of a classifier trained on it is close to that of a classifier trained on the full training data. Recent studies have explored this problem in the context of training neural networks, as seen in works like Toneva et al. (2019); Coleman et al. (2020); Paul et al. (2021); Sorscher et al. (2022); Mindermann et al. (2022). In such scenarios, the oracle can evaluate an underlying loss function exactly, avoiding the need for a noisy estimator.
Our contributions can be summarized as follows:
@@
$\bullet$ We outline a set of conditions under which a non-asymptotic convergence rate of order $O(1/n)$ holds, where $n$ represents the number of iterations of the algorithm, i.e., the number of unlabeled points presented to the algorithm. These conditions enable us to establish convergence rate results for loss-based sampling in linearly separable datasets, with the loss function taking on various forms such as the squared hinge loss function, generalized hinge loss function, or satisfying other specified conditions. Our results provide bounds for both expected loss and the number of sampled points, encompassing different loss-based strategies.
$\bullet$ The results discussed above for loss-based strategies, concerning linear classifiers and linearly separable datasets, rely on a key lemma that may be of independent interest. The conditions of the lemma can be verified for various loss and uncertainty-based acquisition strategies. The convergence rate established in Raj and Bach (2022) for an uncertainty-based strategy follows from this lemma, with a specific choice of the loss function used for training and the loss function used for evaluation. The lemma is derived by generalizing an argument used in the proof of Theorem 3.1 in Raj and Bach (2022) to accommodate different acquisition strategies and loss functions.
@@
$\bullet$ We introduce a framework for establishing convergence rate bounds when sampling points according to an arbitrary, increasing function of the conditional expected loss of a point. This framework is particularly noteworthy as it extends the applicability to a broader set of loss functions compared to our initial conditions. Leveraging the observation that the algorithm behaves akin to a stochastic gradient descent algorithm with an underlying objective function, as referred to as an ”equivalent loss” in Liu and Li (2023), allows us to apply known convergence rate results for stochastic gradient descent algorithms. We present a condition under which explicit bounds on the convergence rate can be established in terms of the sampling probability function, and we provide conditions for this to hold for smooth convex loss functions.
@@
$\bullet$ We propose a learning algorithm that combines a sampling-based acquisition strategy with an adaptive step-size stochastic gradient descent update, achieving the stochastic Polyak’s step size update Loizou et al. (2021) in expectation. We establish a condition on the sampling strategy under which a non-asymptotic convergence rate of order $O(1/n)$ holds for smooth convex loss functions. We present uncertainty and loss-based strategies that satisfy this condition for binary classification, as well as an uncertainty strategy for multi-class classification.
$\bullet$ We present numerical results that demonstrate the efficiency of sampling with stochastic Polyak’s step size and its robustness to loss estimation noise.
@@ 1.1 Related Work
Uncertainty sampling has been utilized for classification tasks since as early as (Lewis and Gale, 1994), and subsequently in many other works, such as (Schohn and Cohn, 2000; Zhu et al., 2010; Yang et al., 2015; Yang and Loog, 2016; Lughofer and Pratama, 2018). Mussmann and Liang (2018) demonstrated that threshold-based uncertainty sampling on a convex loss can be interpreted as performing a pre-conditioned stochastic gradient step on the population zero-one loss. However, none of these works have provided theoretical convergence guarantees. Different variants of uncertainty sampling include margin of confidence sampling, least confidence sampling, and entropy-based sampling (Nguyen et al., 2022). The convergence of margin of confidence sampling was recently studied by Raj and Bach (2022), who showed linear convergence for the hinge loss function for a family of selection probability functions and an algorithm that performs a stochastic gradient descent update with respect to the squared hinge loss function.
A loss-based active learning algorithm was proposed by Yoo and Kweon (2019), comprising a loss prediction module and a target prediction model. The algorithm uses the loss prediction module to compute a loss estimate and prioritizes sampling points with a high estimated loss under the current prediction model. Lahlou et al. (2022) generalize this idea within a framework for uncertainty prediction. However, neither Yoo and Kweon (2019) nor Lahlou et al. (2022) provided theoretical guarantees for convergence rates.
@@
Recent analysis of convergence for loss and uncertainty-based active learning strategies has been presented by Liu and Li (2023). Specifically, they introduced the concept of an equivalent loss, demonstrating that a gradient descent algorithm employing point sampling can be viewed as a stochastic gradient descent algorithm optimizing an equivalent loss function. While they focused on discussing specific cases like sampling proportional to conditional expected loss for loss-based acquisition strategies, we introduce a framework that accommodates sampling based on any continuous increasing function of expected conditional loss. Our framework provides explicit convergence rate bounds in terms of the underlying sampling probability function.
@@
In addition, Loizou et al. (2021) introduced a stochastic gradient descent algorithm featuring an adaptive stochastic Polyak’s step size, which has theoretical convergence guarantees under various assumptions. This algorithm showcased robust performance in comparison to state-of-the-art optimization methods, especially when training over-parametrized models. Our work proposes a novel sampling method that employs stochastic Polyak’s step size in expectation, offering a convergence rate guarantee for smooth convex loss functions, contingent on a condition related to the sampling probability function. Notably, we demonstrate the fulfillment of this condition for logistic regression and binary cross-entropy loss functions, encompassing both a loss-based strategy involving proportional sampling to absolute error loss and an uncertainty sampling strategy. Furthermore, we extend this condition to hold for an uncertainty sampling strategy designed for multi-class classification.
2 Background
Algorithm
@@ 2 Background
We consider the projected stochastic gradient descent algorithm defined as: given an initial value $\theta_{1}$,
@@
where $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta\cap\Theta_{0}}||u-v||$. Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection.
For the choice of the stochastic step size, we consider two cases: (a) Bernoulli sampling with a constant step size, and (b) stochastic Polyak’s step size. For case (a), $z_{t}$ is the product of a constant step size $\gamma$ and a Bernoulli random variable with mean $\pi(x_{t},y_{t},\theta_{t})$. For case (b), $\zeta(x,y,\theta)$ is the ”stochastic” Polyak’s step size, and $z_{t}$ is equal to $\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and is equal to 00 otherwise. Note that using the notation $\pi(x,y,\theta)$ allows for the case when the sampling probability does not depend on the value of the label $y$.
Linearly Separable Data
For binary classification tasks, $\mathcal{Y}=\{-1,1\}$. We say that data is separable if, for every point $(x,y)\in\mathcal{X}\times\mathcal{Y}$, either $y=1$ with probability $1$ or $y=-1$ with probability $1$. The data is linearly separable if there exists $\theta^{*}\in\Theta$ such that $y=\mathrm{sgn}(x^{\top}\theta^{*})$ for every $x\in\mathcal{X}$. Linearly separable data has a $\rho^{*}$-margin if $|x^{\top}\theta^{*}|\geq\rho^{*}$ for every $x\in\mathcal{X}$, for some $\theta^{*}\in\Theta$.
Linear Classifiers
Some of our results are for linear classifiers, where the predicted label of a point $x$ is a function of $x^{\top}\theta$. For example, a model with a predicted label $\mathrm{sgn}(x^{\top}\theta)$ is a linear classifier. For logistic regression, the predicted label is $1$ with probability $\sigma(x^{\top}\theta)$ and $-1$ otherwise, where $\sigma$ is the logistic function defined as $\sigma(z)=1/(1+e^{-z})$.
Smooth Loss Functions
@@ 3 Convergence Rate Guarantees
In this section, we present conditions on the stochastic step size of algorithm (1) under which we can bound the total expected loss and, in some cases, bound the expected number of samples. For Bernoulli sampling with a constant step size, we first provide conditions that allow us to derive bounds for linear classifiers and linearly separable data (Section 3.1). Next, we provide bounds for loss-based sampling, leveraging convergence results known to hold for stochastic gradient descent algorithms (Section 3.2). For sampling combined with stochastic Polyak’s step size, we offer a condition that allows us to establish convergence bounds for loss and uncertainty-based sampling (Section 3.1).
3.1 Bernoulli Sampling with Constant Step Size: First Set of Convergence Conditions
The convergence rate results presented in this section are established using Lemma A.2, which, for space reasons, is provided in Appendix A.1. These results pertain to linear classifiers and linearly separable datasets. We initially address binary classification and subsequently extend our considerations to multi-class classification. The conditions for convergence involve both the loss function $\ell$ utilized by algorithm (1) and a separate loss function $\tilde{\ell}$ employed for evaluating the algorithm’s performance.
Binary Classification
We consider binary classification based on the sign of $x^{\top}\theta$. With a slight abuse of notation, let $\ell(x,y,\theta)\equiv\ell(u)$, ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u)$, and $\pi(x,y,\theta)\equiv\pi(u)$ where $u=yx^{\top}\theta$. We assume that the domain $\mathcal{X}$ is bounded, i.e., there exists $R$ such that $||x||\leq R$ for all $x\in\mathcal{X}$, $||\theta_{1}-\theta^{*}||\leq S$, and that the data is $\rho^{*}$-margin linearly separable (see Section 2 for the definition).
Under prevailing assumptions, the following conditions imply conditions of Lemma A.2: for all $u\in\mathbb{R}$,
$\pi(u)\ell^{\prime}(u)^{2}R^{2}\leq\alpha\tilde{\ell}(u)$ (2)
and
$\pi(u)(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\tilde{\ell}(u).$ (3)
We demonstrate convergence results for a class of loss functions and sampling proportional to either zero-one or absolute error loss, and then for the squared hinge loss function and a smoothed hinge function (which accommodates the hinge loss function as a limit case).
For any $x$, $y$, and $\theta$, the zero-one loss and absolute error loss are defined as $1_{\{y\neq\mathrm{sgn}(x^{\top}\theta)\}}=\ell_{01}(u):=1_{\{u\leq 0\}}$ and $|y-\mathrm{sgn}(x^{\top}\theta)|=\ell_{\mathrm{abs}}(u):=2\left(1_{\{u<0\}}+\frac{1}{2}1_{\{u=0\}}\right)$, respectively, where $u=yx^{\top}\theta$.
We first consider sampling proportional to either zero-one loss or absolute error loss when the loss function satisfies the following conditions:
Assumption 3.1.
In particular, conditions (2) and (3) hold for the case when the loss function $\ell$ is the hinge loss function. For this case, the sampling probability equal to zero-one loss makes algorithm (1) correspond to the well-known perceptron algorithm. This establishes a precise relation between loss-based sampling and the perceptron algorithm.
Theorem 3.2.
Assume that the loss function $\ell$ satisfies Assumption 3.1 and $\rho^{*}>0$. Then, under sampling proportional to zero-one loss with a constant factor $\omega\in(0,1]$, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=c_{1}\rho^{*}/(c_{2}^{2}R^{2})$,
Furthermore, under sampling proportional to absolute error loss with a factor $\omega\in(0,1/2]$ and $\gamma=2c_{1}\rho^{*}/(c_{2}^{2}R^{2})$, the bound in (4) holds but with an additional factor of $2$.
We next consider sampling for the squared hinge loss function $\ell(u)=(1/2)\max\{1-u,0\}^{2}$.

### v3

@@ Abstract
We investigate the convergence rates and data sample sizes required for training a machine learning model using a stochastic gradient descent (SGD) algorithm, where data points are sampled based on either their loss value or uncertainty value. These training methods are particularly relevant for active learning and data subset selection problems. For SGD with a constant step size update, we present convergence results for linear classifiers and linearly separable datasets using squared hinge loss and similar training loss functions. Additionally, we extend our analysis to more general classifiers and datasets, considering a wide range of loss-based sampling strategies and smooth convex training loss functions. We propose a novel algorithm called Adaptive-Weight Sampling (AWS) that utilizes SGD with an adaptive step size that achieves stochastic Polyak’s step size in expectation. We establish convergence rate results for AWS for smooth convex training loss functions. Our numerical experiments demonstrate the efficiency of AWS on various datasets by using either exact or estimated loss values.
@@ 1 Introduction
In practice, when training machine learning models for prediction tasks (classification or regression), one often has access to an abundance of unlabeled data, while obtaining the corresponding labels may entail high costs. This may especially be the case in fields like computer vision, natural language processing, and speech recognition.
Active learning algorithms are designed to efficiently learn a prediction model by employing a label acquisition, with the goal of minimizing the number of labels used to train an accurate prediction model.
@@
Various label acquisition strategies have been proposed, each aiming to select informative points for the underlying model training task; including query-by-committee (Seung et al., 1992), expected model change (Settles et al., 2007), expected error reduction (Roy and McCallum, 2001), expected variance reduction (Wang et al., 2016), and mutual information maximization (Kirsch et al., 2019; Kirsch and Gal, 2022).
@@
A common label acquisition strategy involves estimating uncertainty, which can be viewed as self-disagreement about predictions made by a given model. Algorithms using an uncertainty acquisition strategy are referred to as uncertainty-based active learning algorithms. Different variants of uncertainty strategies include margin of confidence, least confidence, and entropy-based sampling (Nguyen et al., 2022). Recently, a loss-based active learning approach gained attention in research Yoo and Kweon (2019), Lahlou et al. (2022), Nguyen et al. (2021), Luo et al. (2021), and is now applied at scale in industry, such as for training integrity violation classifiers at Meta. This method involves selecting points for which there is a disagreement between the predicted label and the true label, as measured by a loss function. Since the true loss of a data point is unknown prior to the acquisition of the label, in practice, it is estimated using supervised learning. Loss-based sampling aligns with the spirit of the perceptron algorithm (Rosenblatt, 1958), which updates the model only for falsely-classified points.
@@
Convergence guarantees for some uncertainty-based active learning algorithms have recently been established, such as for margin of confidence sampling (Raj and Bach, 2022). By contrast, there are only limited results on the convergence properties of loss-based active learning algorithms, as these only recently been started to be studied, e.g., Liu and Li (2023).
@@
The primary focus of this paper is to establish convergence guarantees for stochastic gradient descent (SGD) algorithms where points are sampled based on their loss. Our work provides new results on conditions that ensure certain convergence rates and bounds on the expected sample size, accommodating various data sampling strategies. Our theoretical results are under assumption that the active learner has access to an oracle that provides unbiased estimate of the conditional expected loss for a point, given the feature vector of the point and the current prediction model. In practice, the loss cannot be evaluated at acquisition time since labels are yet unknown. Instead, a separate prediction model is used for loss estimation. In our experiments, we assess the impact of the bias and noise in such a loss estimator. Our convergence rate analysis accommodates also uncertainty-based data selection, for which we provide new results.
@@
Uncertainty and loss-based acquisition strategies are also of interest for the data subset selection problem, often referred to as core-set selection or data pruning. This problem involves finding a small subset of training data such that the predictive performance of a classifier trained on it is close to that of a classifier trained on the full training data. Recent studies have explored this problem in the context of training neural networks, as seen in works like Toneva et al. (2019); Coleman et al. (2020); Paul et al. (2021); Sorscher et al. (2022); Mindermann et al. (2022). In such scenarios, the oracle can evaluate an underlying loss function exactly, avoiding the need for using a loss estimator.
There is a large body of work on convergence of SGD algorithms, e.g. see Bubeck (2015); Nesterov (2018). These results are established for SGD algorithms under either constant, diminishing or adaptive step sizes. Recently, Loizou et al. (2021), studied SGD with the stochastic Polyak’s step size, depending on the ratio of the loss and the squared gradient of the loss of a point. Our work is different as we consider convergence of SGD under sampling of points.
1.1 Summary of our Contributions
Our contributions can be summarizes as given in the following points:
@@ 1.1 Summary of our Contributions
$\bullet$ For SGD with a constant step size, we present conditions under which a non-asymptotic convergence rate of order $O(1/n)$ holds, where $n$ represents the number of iterations of the algorithm, i.e., the number of unlabeled points presented to the algorithm. These conditions enable us to establish convergence rate results for loss-based sampling in the case of linear classifiers and linearly separable datasets, with the loss function taking on various forms such as the squared hinge loss function, generalized hinge loss function, or satisfying other specified conditions. Our results provide bounds for both expected loss and the number of sampled points, encompassing different loss-based strategies. These results are established by using a convergence rate lemma that may be of independent interest.
@@
$\bullet$ For SGD with a constant step size, we provide new convergence rate results for more general classifiers and datasets, with sampling of points according to an increasing function $\pi$ of the conditional expected loss of a point. In this case, we present conditions for smooth convex training loss functions under which a non-asymptotic convergence rate of order $O(\Pi^{-1}(1/\sqrt{n}))$ holds, where $\Pi$ is the primitive function of $\pi$. These results are established by leveraging the fact that the algorithm behaves akin to a SGD algorithm with an underlying objective function, as referred to as an equivalent loss in Liu and Li (2023), allowing us to apply known convergence rate results for SGD algorithms.
@@
$\bullet$ We propose Adaptive-Weight Sampling (AWS), a novel learning algorithm that combines a sampling-based acquisition strategy with an adaptive step-size SGD update, achieving the stochastic Polyak’s step size update in expectation, which can be used with any differentiable loss function. We establish a condition under which a non-asymptotic convergence rate of order $O(1/n)$ holds for AWS with smooth convex loss functions. We present uncertainty and loss-based strategies that satisfy this condition for binary classification, as well as an uncertainty strategy for multi-class classification.
$\bullet$ We present numerical results that demonstrate the efficiency of AWS on various datasets.
@@ 1.2 Related Work
Uncertainty sampling has been utilized for classification tasks since as early as (Lewis and Gale, 1994), and subsequently in many other works, such as (Schohn and Cohn, 2000; Zhu et al., 2010; Yang et al., 2015; Yang and Loog, 2016; Lughofer and Pratama, 2018). Mussmann and Liang (2018) demonstrated that threshold-based uncertainty sampling on a convex loss can be interpreted as performing a pre-conditioned stochastic gradient step on the population zero-one loss. However, none of these works have provided theoretical convergence guarantees.
The convergence of margin of confidence sampling was recently studied by Raj and Bach (2022), who demonstrated linear convergence for linear classifiers and linearly separable datasets, specifically for the hinge loss function, for a family of selection probability functions and an algorithm that performs a SGD update with respect to the squared hinge loss function. However, our results for linear classifiers and linearly separable datasets differ, as our focus lies on loss-based sampling strategies and providing bounds on the convergence rate of a loss function and the expected number of sampled points. These results are established using a convergence rate lemma, which may be of independent interest. It is noteworthy that the convergence rate for uncertainty-based sampling, as in Theorem 3.1 of Raj and Bach (2022), can be derived by checking the conditions of the convergence rate lemma.
@@
A loss-based active learning algorithm was proposed by Yoo and Kweon (2019), comprising a loss prediction module and a target prediction model. The algorithm uses the loss prediction module to compute a loss estimate and prioritizes sampling points with a high estimated loss under the current prediction model. Lahlou et al. (2022) generalize this idea within a framework for uncertainty prediction. However, neither Yoo and Kweon (2019) nor Lahlou et al. (2022) provided theoretical guarantees for convergence rates. Recent analysis of convergence for loss and uncertainty-based active learning strategies has been presented by Liu and Li (2023). Specifically, they introduced the concept of an equivalent loss, demonstrating that a gradient descent algorithm employing point sampling can be viewed as a SGD algorithm optimizing an equivalent loss function. While they focused on specific cases like sampling proportional to conditional expected loss, our results allow for sampling based on any continuous increasing function of expected conditional loss, and provide explicit convergence rate bounds in terms of the underlying sampling probability function.
@@
In addition, Loizou et al. (2021) introduced a SGD algorithm featuring an adaptive stochastic Polyak’s step size, which has theoretical convergence guarantees under various assumptions. This algorithm showcased robust performance in comparison to state-of-the-art optimization methods, especially when training over-parametrized models. Our work proposes a novel sampling method that employs stochastic Polyak’s step size in expectation, offering a convergence rate guarantee for smooth convex loss functions, contingent on a condition related to the sampling probability function. Notably, we demonstrate the fulfillment of this condition for logistic regression and binary cross-entropy loss functions, encompassing both a loss-based strategy involving proportional sampling to absolute error loss and an uncertainty sampling strategy. Furthermore, we extend this condition to hold for an uncertainty sampling strategy designed for multi-class classification.
2 Problem Statement
@@
We consider the projected SGD algorithm defined as: given an initial value $\theta_{1}\in\Theta$,
@@
where $\ell:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\mathbb{R}$ is a training loss function, $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta_{0}}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta_{0}}||u-v||$. Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection. For binary classification tasks, we assume $\mathcal{Y}=\{-,1,1\}$. For every $t>0$, we define $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
For the choice of the stochastic step size, we consider two cases: (a) Constant-Weight Sampling: a Bernoulli sampling with a constant step size, and (b) Adaptive-Weight Sampling: a sampling that achieves stochastic Polyak’s step size in expectation. For case (a), $z_{t}$ is the product of a constant step size $\gamma$ and a Bernoulli random variable with mean $\pi(x_{t},y_{t},\theta_{t})$. For case (b), $\zeta(x,y,\theta)$ is the "stochastic" Polyak’s step size, and $z_{t}$ is equal to $\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and is equal to 00 otherwise. Note that using the notation $\pi(x,y,\theta)$ allows for the case when the sampling probability does not depend on the value of the label $y$.
For a loss-based sampling, $\pi$ is an increasing function of some loss function $\ell^{\star}$, which does not necessarily correspond to the training loss function $\ell$. Specifically, for a binary classifier with $p(x,y,\theta)$ denoting the expected prediction label, sampling proportional to the absolute error loss is defined as $\pi(\ell^{*})=\omega\ell^{*}$ where $\ell^{*}(x,y,\theta)=|y-p(x,y,\theta)|$ and $\omega\in(0,1/2]$. For an uncertainty-based sampling, $\pi$ is a function of some quantity reflecting the uncertainty of the prediction model.
Our focus is on finding convergence conditions for algorithm (1) and convergence rates under these conditions, as well as bounds on the expected number of points sampled by the algorithm.
Additional Assumptions and Notation
For binary classification, we say that data is separable if, for every point $(x,y)\in\mathcal{X}\times\mathcal{Y}$, either $y=1$ with probability $1$ or $y=-1$ with probability $1$. The data is linearly separable if there exists $\theta^{*}\in\Theta$ such that $y=\mathrm{sgn}(x^{\top}\theta^{*})$ for every $x\in\mathcal{X}$. Linearly separable data has a $\rho^{*}$-margin if $|x^{\top}\theta^{*}|\geq\rho^{*}$ for every $x\in\mathcal{X}$, for some $\theta^{*}\in\Theta$.
Some of our results are for linear classifiers, where the predicted label of a point $x$ is a function of $x^{\top}\theta$. For example, a model with a predicted label $\mathrm{sgn}(x^{\top}\theta)$ is a linear classifier. For logistic regression, the predicted label is $1$ with probability $\sigma(x^{\top}\theta)$ and $-1$ otherwise, where $\sigma$ is the logistic function defined as $\sigma(z)=1/(1+e^{-z})$. For binary classification, we model the prediction probability of the positive label as $\sigma(x^{\top}\theta)$, where $\sigma:\mathbb{R}_{+}\rightarrow[0,1]$ is an increasing function, and $\sigma(-u)+\sigma(u)=1$ for all $u\in\mathbb{R}$. The absolute error loss takes value $1-\sigma(x^{\top}\theta)$ if $y=1$ or value $\sigma(x^{\top}\theta)$ if $y=-1$, which corresponds to $1-\sigma(yx^{\top}\theta)$. The binary cross-entropy loss for a point $(x,y)$ under model parameter $\theta$ can be written as $\ell(x,y,\theta)=-\log(\sigma(yx^{\top}\theta))$. Hence, absolute error loss-based sampling corresponds to the sampling probability function $\pi(\ell)=1-e^{-\ell}$.
@@ 3 Convergence Rate Guarantees
In this section, we present conditions on the stochastic step size of algorithm (1) under which we can bound the total expected loss and the expected number of samples. For the Constant-Weight Sampling, we provide conditions that allow us to derive bounds for linear classifiers and linearly separable datasets and more general cases. For Adaptive-Weight Sampling, we offer a condition that allows us to establish convergence bounds for both loss and uncertainty-based sampling.
3.1 Constant-Weight Sampling
Linear Classifiers and Linearly Separable Datasets
We focus on binary classification and briefly discuss extension to multi-class classification. We consider the linear classifier with the predicted label $\hbox{sgn}(x^{\top}\theta)$. With a slight abuse of notation, let $\ell(x,y,\theta)\equiv\ell(u)$ and $\pi(x,y,\theta)\equiv\pi(u)$ where $u=yx^{\top}\theta$. We assume that the domain $\mathcal{X}$ is bounded, i.e., there exists $R$ such that $||x||\leq R$ for all $x\in\mathcal{X}$, $||\theta_{1}-\theta^{*}||\leq S$ for some $S\geq 0$, and that the data is $\rho^{*}$-margin linearly separable.
We present convergence rate results for the training loss function corresponding to the squared hinge loss function, i.e. $\ell(u)=(1/2)\max\{1-u,0\}^{2}$. Our additional results also cover other cases, including a class of smooth convex loss functions and a generalized smooth hinge loss function, which are presented in the Appendix.